# CT-based Anatomical Segmentation for Thoracic Surgical Planning: A Benchmark Study for 3D U-shaped Deep Learning Models

### 05 Feb 2024 v1

@@ Abstract
Recent rising interests in patient-specific thoracic surgical planning and simulation require efficient and robust creation of digital anatomical models from automatic medical image segmentation algorithms. Deep learning (DL) is now state-of-the-art in various radiological tasks, and U-shaped DL models have particularly excelled in medical image segmentation since the inception of the 2D UNet. To date, many variants of U-shaped models have been proposed by the integration of different attention mechanisms and network configurations. Leveraging the recent development of large multi-label databases, systematic benchmark studies for these models can provide valuable insights for clinical deployment and future model designs, but such studies are still rare. We conduct the first benchmark study for variants of 3D U-shaped models (3DUNet, STUNet, AttentionUNet, SwinUNETR, FocalSegNet, and a novel 3D SwinUnet with four variants) with a focus on CT-based anatomical segmentation for thoracic surgery. Our study systematically examines the impact of different attention mechanisms, number of resolution stages, and network configurations on segmentation accuracy and computational complexity. To allow cross-reference with other recent benchmarking studies, we also included a performance assessment of the BTCV abdominal structural segmentation. With the STUNet ranking at the top, our study demonstrated the value of CNN-based U-shaped models for the investigated tasks and the benefit of residual blocks in network configuration designs to boost segmentation performance.
@@ I Introduction
Deep learning (DL) approaches, such as convolutional neural networks (CNNs) have dominated the state-of-the-art performance in various radiological tasks. With their quick inference time, they offer a tool to enable efficient digital twin construction for thoracic surgical planning, simulation, and intra-operative monitoring. U-shaped models, pioneered by the 2D UNet [7], stand out among DL segmentation models [8] for their robust performance and elegant architecture. The typical U-shaped architecture comprises three key elements: an encoder for learning relevant image features and compressing them into lower-dimensional embeddings, a decoder for expanding these embeddings and producing the final segmentation, and skip connections that maintain fine-grained details during upsampling by aggregating feature maps across encoder and decoder layers. Since the first inception, major efforts have been dedicated to adapting the 2D framework to 3D [9], exploring different backbones for the encoder/decoder [10, 11, 12, 13, 14, 15], updating the resolution levels [16], and experimenting with novel network configurations [17]. To better understand the impact of these enhancements and investigate the application of 3D anatomical reconstruction for thoracic surgical planning and simulation, a comprehensive benchmark study would be highly instrumental, but has yet to be conducted.
@@ I-C Novelty and Contributions
  • Benchmark different U-shaped models in 3D image segmentation for thoracic surgery and BTCV challenge anatomies by considering their accuracy and computational complexity
  • Investigate the impacts of different network configuration designs, including the number of resolution stages (number of upsampling/downsampling operations) for all selected models, and different operations for skip connections, downsampling, and upsampling in a pure Swin Transformer-based U-shaped model
@@ IV Discussion
Our presented study primarily focuses on the impacts of attention mechanisms, number of resolution stages, and network configuration designs for U-shaped deep learning models. As a result, we selected the most popular and representative models for the themes of our investigation, instead of benchmarking an exhaustive list of U-shaped models. For the model performance, it is possible that the insights drawn from the experiments may be task- and modality-specific [38, 50], and should be verified further with additional benchmarking datasets.

### 14 Mar 2024 v2

# Architecture Analysis and Benchmarking of 3D U-shaped Deep Learning Models for Thoracic Anatomical Segmentation

@@ Abstract
Recent rising interests in patient-specific thoracic surgical planning and simulation require efficient and robust creation of digital anatomical models from automatic medical image segmentation algorithms. Deep learning (DL) is now state-of-the-art in various radiological tasks, and U-shaped DL models have particularly excelled in medical image segmentation since the inception of the 2D UNet. To date, many variants of U-shaped models have been proposed by the integration of different attention mechanisms and network configurations. Systematic benchmark studies which analyze the architecture of these models by leveraging the recent development of the multi-label databases, can provide valuable insights for clinical deployment and future model designs, but such studies are still rare. We conduct the first systematic benchmark study for variants of 3D U-shaped models (3DUNet, STUNet, AttentionUNet, SwinUNETR, FocalSegNet, and a novel 3D SwinUnet with four variants) with a focus on CT-based anatomical segmentation for thoracic surgery. Our study systematically examines the impact of different attention mechanisms, the number of resolution stages, and network configurations on segmentation accuracy and computational complexity. To allow cross-reference with other recent benchmarking studies, we also included a performance assessment of the BTCV abdominal structural segmentation. With the STUNet ranking at the top, our study demonstrated the value of CNN-based U-shaped models for the investigated tasks and the benefit of residual blocks in network configuration designs to boost segmentation performance.
@@ I Introduction
Deep learning (DL) approaches, such as convolutional neural networks (CNNs) have dominated the state-of-the-art performance in various radiological tasks. With their quick inference time, they offer a tool to enable efficient digital twin construction for thoracic surgical planning, simulation, and intra-operative monitoring. U-shaped models, pioneered by the 2D UNet [7], stand out among DL segmentation models [8] for their robust performance and elegant architecture. The typical U-shaped architecture comprises three key elements: an encoder for learning relevant image features and compressing them into lower-dimensional embeddings, a decoder for expanding these embeddings and producing the final segmentation, and skip connections that maintain fine-grained details during upsampling by aggregating feature maps across encoder and decoder layers. Since the first inception, major efforts have been dedicated to adapting the 2D framework to 3D [9], exploring different backbones for the encoder/decoder [10, 11, 12, 13, 14, 15], updating the resolution stages [16], and experimenting with novel network configurations [17]. To better understand the impact of these enhancements and investigate the application of 3D anatomical reconstruction for thoracic surgical planning and simulation, a comprehensive benchmark study, providing the models’ architectural characteristics, would be highly instrumental, but has yet to be conducted.
@@ I-C Novelty and Contributions
  • Benchmark state-of-the-art U-shaped models, including our novel 3D SwinUNet and its four variants in 3D image segmentation for thoracic surgery and BTCV challenge anatomies by considering their accuracy and computational complexity
  • Investigate the impacts of different network configuration designs, including the number of resolution stages (number of upsampling/downsampling operations) and different operations for skip connections, downsampling, and upsampling in a pure Swin Transformer-based U-shaped model
@@ IV Discussion
Our presented study primarily focuses on the impacts of attention mechanisms, number of resolution stages, and network configuration designs for U-shaped deep learning models. As a result, we selected the most popular and representative models for the themes of our investigation, instead of analyzing an exhaustive list of U-shaped models. For the model performance, it is possible that the insights drawn from the experiments may be task- and modality-specific [38, 50], and should be verified further with additional benchmarking datasets.